18.1 Mastery #3 Q2
Solubility to Ksp

A student measures the molar solubility of iron(II) hydroxide in a water solution to be 1.29×10-5 M.  
Based on her data, the solubility product constant for this compound is ___.
Answer #1

At equilibrium:

Fe(OH)2 <----> Fe2+ + 2 OH-

   s 2s

Ksp = [Fe2+][OH-]^2

Ksp = (s)*(2s)^2

Ksp = 4(s)^3

Ksp = 4(1.29*10^-5)^3

Ksp = 8.587*10^-15

Answer: 8.59*10^-15